# LaV-CoT: Language-Aware Visual CoT with Multi-Aspect Reward Optimization for Real-World Multilingual VQA

### v3

@@ Abstract.
[…]
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Companion Proceedings of the ACM Web Conference 2026; April 13–17, 2026; Dubai, UAE**footnotetext: This work is a preprint.
@@ 1. Introduction
Large Vision Language Models (VLMs) (Li et al., 2022; Alayrac et al., 2022; Liu et al., 2023; Lu et al., 2024; Wang et al., 2024a, b; Li et al., 2024b; Sun et al., 2024; Bai et al., 2025; Zhu et al., 2025; Wang et al., 2025) have advanced rapidly in recent years, demonstrating impressive performance on multimodal tasks such as image captioning (Wang et al., 2023), visual question answering (VQA) (Chen et al., 2022b), and complex reasoning (Li et al., 2024a). By integrating visual and textual modalities, these models learn rich semantic representations and enable a wide range of real-world applications. With the growing demand for global accessibility, multilingual visual question answering (mVQA) (Changpinyo et al., 2023) has emerged as a critical capability for deploying VLMs at scale.
@@
Despite recent progress in vision language models (VLMs), achieving reliable reasoning in multilingual multimodal settings remains challenging. Textual Chain-of-Thought (CoT) (Wei et al., 2022) reasoning improves interpretability and supports step-by-step inference (Wei et al., 2022). However, when applied to multilingual visual question answering (mVQA), a text-only rationale often under-exploits visual cues and weakens visual grounding. Orthogonal to the choice of CoT, current VLMs for mVQA still exhibit persistent issues: i) Limited Interpretability. Direct answers produced by recent VLMs (Bai et al., 2025; Wang et al., 2025) remain as black box, hindering model transparency and interpretability. (ii) Visual–textual misalignment. Textual CoT approaches (Wei et al., 2022; Xu et al., 2025) introduce intermediate natural language reasoning steps to enhance performance. However, these text-only rationales often under-exploit visual cues, leading to weak visual grounding. (iii) Language inconsistency. Recent Visual CoT methods (Shao et al., 2024a; Zhao et al., 2025) incorporate visual cues to enable multimodal reasoning, but still suffer from inconsistency across languages. Such limitations constrain the deployment of VLMs in international products, which demand both robustness and interpretable reasoning.
To address these challenges, we propose LaV-CoT, a Language-aware Visual CoT reasoning framework with Multi-Aspect Reward Optimization. As illustrated in Figure 1, LaV-CoT introduces an interpretable multi-stage reasoning pipeline that integrates four key components: (1) Text Summary with Bounding Box (BBox) (Yang et al., 2022), (2) Language Identification, (3) Spatial Object-level Captioning (Huang, 2024), and (4) Step-by-step Logical Reasoning (Surís et al., 2023; Zhang et al., 2024; Xu et al., 2025). This structured pipeline explicitly disentangles language and visual reasoning, enabling fine-grained cross-modal alignment and improving interpretability in multilingual settings.
@@
Figure 1. Overview of Lav-CoT: (a) Direct model answers may be incorrect and exhibit language inconsistency. (b) Incorporating CoT reasoning enhances reasoning transparency but still exhibits linguistic inconsistency. (c) Our method introduces a multi-stage reasoning pipeline, yielding accurate and consistent final answers.:
A further challenge lies in constructing high-quality multilingual reasoning data. Manual annotation is costly and often infeasible at scale. To this end, we design an automatic data curation method that generates multilingual CoT annotations via iterative generation, correction, and refinement (Zelikman et al., 2022; Lightman et al., 2023; Shao et al., 2025). This scalable process produces structured and verifiable reasoning traces, ensuring that training data captures both linguistic fidelity and multimodal reasoning quality.
@@
On the training side, LaV-CoT adopts a two-stage paradigm combining Supervised Fine-Tuning (SFT) (Li et al., 2022; Chen et al., 2022b; Bai et al., 2025) with Language-aware Group Relative Policy Optimization (GRPO) (Lu et al., 2023; Shao et al., 2024b; Liu et al., 2025). Unlike standard optimization methods, our GRPO variant is guided by verifiable multi-aspect rewards—including language consistency reward, text segments and object count reward, final answer edit distance reward and format reward—yielding stable optimization and robust reasoning across languages and modalities.
[…]
@@
  • We propose LaV-CoT, the first framework to unify language-aware visual CoT reasoning with multi-aspect reward optimization for multilingual multimodal tasks.
@@
  • We develop a language-aware GRPO algorithm with verifiable multi-aspect rewards, enhancing reasoning robustness and cross-modal alignment.
  • We validate our framework through extensive experiments on public benchmarks and further reinforce its effectiveness via a online A/B test, demonstrating state-of-the-art multilingual multimodal reasoning performance.
Figure 2. The framework includes an automated data generation pipeline, which leverages a multi-step reasoning process comprising (a) Text Summary with BBox, (b) Language Identification, (c) Spatial Object-level Image Captioning, and (d) Step-by-step Logical Reasoning. This reasoning pipeline first produces vanilla CoT annotations, which are then iteratively refined through rigorous verification to ensure high-quality supervision. The generated data subsequently supports a two-stage training paradigm combining SFT with GRPO. Reward computation considers language consistency, count accuracy, and the edit distance between predicted and ground-truth answers, collectively enhancing structural understanding and reinforcing robust reasoning capabilities.:
@@ 2. RELATED WORKS
2.1. Large Multilingual Vision-Language Reasoning Model
Large multilingual vision-language reasoning models bridge complex visual understanding and cross-lingual semantic reasoning. Early works focused on combining visual features with symbolic reasoning to address structured visual tasks (Andreas et al., 2016; Liang et al., 2018). With the advent of large language models (LLMs), recent approaches leverage their powerful reasoning abilities to interpret and generate multimodal content across multiple languages(Alayrac et al., 2022; Chen et al., 2022b). To address multilingual challenges, models now incorporate text encoders and alignment strategies, enabling seamless reasoning across diverse languages (Karoui et al., 2023; Zhang et al., 2023). Visual representation advances, such as cognition-focused visual tokens, boost fine-grained reasoning (Bigverdi et al., 2025; Chen et al., 2025). Innovative frameworks treat the language model as a reasoning agent interacting dynamically with specialized visual modules, supporting flexible and interpretable workflows, as in VISPROG (Asai et al., 2022). And (Ahmat et al., 2025) propose a Multi-grained Multilingual Vision-Language Pre-training (M2-VLP) model, which uses a unified multi-grained contrastive learning paradigm to integrate cross-language and cross-modal alignment.(Ranaldi et al., 2025) proposed the R2-MultiOmnia method, which guides key elements in the model’s abstract reasoning process and optimizes the reasoning trajectory through self-correction, enabling Multimodal Large Language Models (MLLMs) to maintain consistent reasoning capabilities across different languages and improve modal performance.
2.2. Chain-of-thought in Large Language Models
@@
Chain-of-thought (CoT) prompting (Wei et al., 2022) provides a systematic, stepwise reasoning framework that helps large language models tackle challenging problems such as commonsense reasoning (Wei et al., 2022; Kumar et al., 2022) and logical deduction (Cobbe et al., 2021; Wei et al., 2022). By breaking down complex questions into a sequence of intermediate reasoning steps, CoT guides the model to produce detailed, incremental solutions rather than direct answers. Recent research has shown that incorporating CoT strategies significantly boosts the reasoning abilities of vision-language models (VLMs). For example, Prism (Chen et al., 2022a) separates the reasoning pipeline into distinct perception and inference stages to enhance understanding. MSG (Luo et al., 2023) introduces forced Chain-of-Thoughts, laying the groundwork for structured prompting methods. Approaches like Distilling CoT (Hsieh et al., 2023) and Visual Program Distillation (Hu et al., 2024) focus on transferring CoT reasoning patterns into VLMs. Visual CoT (Shao et al., 2025) enhances interpretability by producing bounding boxes that highlight relevant image regions alongside textual answers, providing clearer visual grounding of the reasoning process. Meanwhile, LLaVA-CoT (Lin et al., 2023) employs a structured Chain-of-Thought mechanism, which systematically decomposes complex reasoning tasks and thereby achieves improved accuracy in multimodal reasoning. Building upon these advancements, our work proposes a systematic and interpretable CoT framework explicitly designed to address the challenges of multilingual vision-language reasoning.
2.3. Reinforcement Tuning
Reinforcement Learning (RL) (Sutton and Barto, 2018; Mnih et al., 2015; Silver et al., 2016) is a fundamental paradigm in machine learning, where an agent learns to interact with an environment by performing actions, receiving feedback in the form of rewards, and iteratively updating its policy to maximize cumulative long-term returns. Classical RL algorithms, such as Q-learning (Watkins and Dayan, 1992), have been successfully applied across diverse domains including robotics, game playing (e.g., AlphaGo), and autonomous control. With the advent of large language models (LLMs), Reinforcement Learning with Human Feedback (RLHF) (Christiano et al., 2017; Stiennon et al., 2020) has become a prominent technique to fine-tune models by leveraging human preference signals. RLHF commonly employs policy optimization algorithms like Proximal Policy Optimization (PPO) (Ouyang et al., 2022) and Direct Preference Optimization (DPO) (Rafailov et al., 2023) to steer model behavior, thereby enhancing alignment, coherence, and helpfulness in generated responses. More recently, advanced policy optimization variants such as Group Relative Policy Optimization (GRPO) (Lu et al., 2023) have been proposed, which modify the standard value-function estimator with a group-relative advantage function computed over multiple candidate responses per prompt. Moreover, RL-based fine-tuning strategies have also been successfully extended to vision-language models (VLMs). V-DPO (Xie et al., 2024) uses vision-guided Direct Preference Optimization to enhance visual context learning at training time. In our work, we design verifiable multi-aspect reward functions to enhance the granularity and reliability of feedback signals during reinforcement learning.
@@ 3. Method
Our framework enables a progressive, step-by-step reasoning process that strengthens the reasoning capabilities of Vision-Language Models (VLMs). Leveraging structured multi-stage thinking, it establishes a systematic and efficient reasoning pipeline. To ensure high-quality training data, the data curation stage incorporates iterative CoT evaluations to identify and correct errors until each reasoning sequence is fully validated. Finally, a two-stage training strategy with a novel, multi-aspect reward design is employed to obtain the final optimized model.
3.1. Multi-stage Reasoning
The multi-stage reasoning design is motivated by the way humans naturally approach image understanding: when examining a document image, we first locate salient text regions and retain their summarized content, then identify the language, recognize objects and their spatial relationships, and finally integrate all information to reason step by step toward an answer. As illustrated in Figure 2, our model performs reasoning through four structured stages:
@@ 3.1. Multi-stage Reasoning
3.2. Dataset Curation
Most existing VQA datasets lack the detailed reasoning processes necessary to effectively train the multilingual reasoning vlm model. To address this limitation, we compile a new dataset by integrating samples from several widely used VQA benchmarks, resulting in a total of 148k image–question–CoT–answer pairs, as illustrated in Table 3.
Specifically, as shown in Algorithm 1, we start from the original question–answer triplets $\langle I,Q,A\rangle$, where $I$ denotes the image, $Q$ the question, and $A$ the corresponding answer. We first prompt a GPT-based generator $f_{\text{gen}}$ to produce an initial sequence of vanilla Chain-of-Thought (CoT) steps $T_{\text{init}}=\{s_{i}\}_{i=1}^{|T_{\text{init}}|}$. Next, we prompt an evaluator $f_{\text{eval}}$ to score each step $s_{i}\in T_{\text{init}}$. For any step whose score falls below the threshold $\tau$, we iteratively perform the following procedure: First, we apply the evaluator to the step and then locate the erroneous part, denoted as $s_{\text{error}}$, using the function $\text{Locate}(f_{\text{eval}}(s_{i}))$. Next, a corrected step $s_{i}$ is generated by applying the function Correct to $s_{\text{error}}$, i.e., $s_{i}$ is updated as $\text{Correct}(f_{\text{gen}}(s_{\text{error}}))$. The corrected step then replaces the original step in the sequence $T_{\text{init}}$. Finally, the updated step is re-evaluated to obtain the new score using $f_{\text{eval}}(s_{i})$. This evaluation–correction–update loop repeats until all steps in $T_{\text{init}}$ exceed the threshold. The final verified Chain-of-Thought sequence is denoted as $T$, and the output dataset consists of $\langle I,Q,T,A\rangle$. Detailed prompt design are shown in Appendix B.
Our dataset covers 13 languages, including English (EN), Chinese (ZH), Portuguese (PT), Arabic (AR), Turkish (TR), Russian (RU), German (DE), French (FR), Italian (IT), Japanese (JA), Korean (KO), Thai (TH), and Vietnamese (VI). These languages represent a diverse range of linguistic families. Our datasets are selected from diverse sources, including COCO2017(Lin et al., 2014), a large-scale dataset for object detection, segmentation, and image captioning containing everyday scenes with common objects in context; Visual Genome(Krishna et al., 2016), which provides densely annotated images linking objects, attributes, relationships, and region descriptions for visual reasoning; GQA(Hudson and Manning, 2019), emphasizing compositional reasoning over real-world images with complex question-answer pairs; OCR-VQA(Mishra et al., 2019) and TextVQA(Singh et al., 2019), both focusing on reading and reasoning over text within images for text-centric visual question answering; Llava-Pretrain(Lin et al., 2023), a multimodal pretraining dataset supporting various VQA and reasoning tasks across multiple domains; and MTVQA(Tang et al., 2024), a multilingual text-based VQA dataset designed to evaluate cross-lingual multimodal understanding.

### v4

# LaV-CoT: Language-Aware Visual CoT with Multi-Aspect Reward Optimization for Multilingual Text-Centric VQA

@@ Abstract.
Multilingual Text-Centric Visual Question Answering (TEC-VQA) has become crucial for real-world applications, as it requires fine-grained understanding and reasoning over multilingual scene text. Recent advances in vision-language models (VLMs) have demonstrated strong potential in tackling multimodal tasks. However, most existing approaches rely primarily on textual Chain-of-Thought (CoT) and provide limited support for multilingual multimodal reasoning. To address this gap, we introduce LaV-CoT, the first Language-aware Visual CoT framework with Multi-Aspect Reward Optimization. LaV-CoT incorporates an interpretable multi-stage reasoning pipeline consisting of text summary with bounding box, language identification, spatial object-level captioning, and step-by-step logical reasoning. To improve reasoning accuracy and cross-lingual generalization, we propose a novel verifiable Multi-Aspect Reward Optimization in addition to supervised fine-tuning that incorporates rewards for linguistic consistency, structural fidelity, and response accuracy. Extensive evaluations on public datasets, including MMMB, Multilingual MMBench, and MTVQA, show that LaV-CoT outperforms open-source models of similar size by up to $\sim$9.5% accuracy, even surpassing open-source models more than twice its size, and further exceeding several state-of-the-art proprietary models. Moreover, LaV-CoT has been integrated into our online Intelligent Document Processing platform. A further online A/B test demonstrates an $\sim$8.7% improvement in acceptance rate, validating its effectiveness in industrial deployment and commercial applications. Our code is available at this repository.
Multilingual Vision-Language Models; Multilingual Text-Centric VQA; Visual Chain-of-Thought Reasoning; Reinforcement Tuning
copyright: acmlicensedjournalyear: 2026copyright: ccconference: Proceedings of the ACM Web Conference 2026; April 13–17, 2026; Dubai, United Arab Emiratesbooktitle: Proceedings of the ACM Web Conference 2026 (WWW ’26), April 13–17, 2026, Dubai, United Arab Emiratesdoi: 10.1145/3774904.3792793isbn: 979-8-4007-2307-0/2026/04ccs: Computing methodologies Machine learningccs: Computing methodologies Computer visionccs: Computing methodologies Natural language processing
@@ 1. Introduction
Large Vision-Language Models (VLMs) (Li et al., 2022; Alayrac et al., 2022; Liu et al., 2023b; Lu et al., 2024; Wang et al., 2024a, b; Li et al., 2024b; Sun et al., 2025; Bai et al., 2025; Zhu et al., 2025; Wang et al., 2025; Li et al., 2025; GLM et al., 2024) have advanced rapidly in recent years, demonstrating strong performance on multimodal tasks such as image captioning (Li et al., 2022; Wang et al., 2023; Zhou et al., 2019), text–region grounding (Zhang et al., 2022; Wan et al., 2024) and visual question answering (Chen et al., 2022; Zhou et al., 2019). Driven by the growing demand from international industrial applications, multilingual Text-Centric Visual Question Answering (TEC-VQA) (Tang et al., 2025; Zeng et al., 2021; Mishra et al., 2019) has emerged as a key task to evaluate models’ ability to interpret and reason over multilingual text in images. Recent VLMs have begun to conduct preliminary evaluations on multilingual TEC-VQA benchmarks.
@@
Although recent general VLMs are trained on an increasing number of VQA datasets, they still suffer from a scarcity of multilingual text-centric data. Expert models (Sun et al., 2025; Li et al., 2024b) include additional relevant datasets but remain limited in multimodal reasoning, leaving multilingual TEC-VQA a persistent challenge. As illustrated in Figure 1, three major issues persist: (i) Limited Interpretability. Direct answers produced by recent VLMs (Bai et al., 2025; Wang et al., 2025) remain as black boxes, hindering model transparency and interpretability. (ii) Visual–textual misalignment. Textual CoT approaches (Wei et al., 2022; Xu et al., 2025) introduce intermediate natural language reasoning steps to enhance performance. However, these text-only rationales often under-exploit visual cues, leading to weak visual grounding. (iii) Language inconsistency. Recent Visual CoT methods (Shao et al., 2024a; Zhao et al., 2025) incorporate visual cues to enable multimodal reasoning, but still suffer from inconsistency across languages.
@@
Figure 1. Overview of LaV-CoT: (a) Direct model answers may be incorrect with limited interpretability. (b) Textual or Visual CoT reasoning improves reasoning transparency but still suffers from weak visual grounding and linguistic inconsistency. (c) Our method proposes an interpretable multi-stage reasoning pipeline and gets the correct final answer.
Reasoning pipeline
To address these challenges, we propose LaV-CoT, a Language-aware Visual CoT reasoning framework with Multi-Aspect Reward Optimization. LaV-CoT introduces an interpretable multi-stage reasoning pipeline that integrates four complementary components: (1) Text Summary with Bounding Box (Yang et al., 2022), which localizes and extracts textual content from the image; (2) Language Identification, which ensures accurate linguistic understanding and processing; (3) Spatial Object-level Captioning (Li et al., 2019), which grounds visual entities within the scene; and (4) Step-by-step Logical Reasoning (Surís et al., 2023; Zhang et al., 2024; Xu et al., 2025), where the model systematically analyzes the question and integrates information across modalities to reach the final answer. Together, these components form a structured and interpretable reasoning process, enabling fine-grained cross-modal alignment and improving interpretability in multilingual multimodal tasks. A further challenge lies in constructing high-quality multilingual reasoning data due to the high cost of manual annotation. To this end, we design an automatic data curation method that generates structured and verifiable multilingual multimodal CoT annotations through iterative generation, correction, and refinement (Zelikman et al., 2022; Lightman et al., 2023; Shao et al., 2024a), thereby preserving both linguistic fidelity and reasoning quality at scale.
@@
Recent studies have shown that reinforcement learning (RL) leads to better generalization in both rule-based and visual reasoning tasks (Chu et al., 2025). However, existing RL approaches are often limited by coarse-grained, binary rewards that fail to capture fine-grained reasoning quality for multilingual TEC-VQA. Therefore, we propose a two-stage paradigm that combines supervised fine-tuning (SFT) (Li et al., 2022; Chen et al., 2022; Bai et al., 2025) with language-aware Group Relative Policy Optimization (GRPO) (Lu et al., 2023; Shao et al., 2024b; Liu et al., 2025a). Unlike standard optimization methods, our GRPO variant is guided by verifiable multi-aspect rewards, including language consistency, text segment and object count accuracy, final answer edit distance, and format compliance. This multi-aspect reward design enables stable optimization and robust reasoning across diverse languages and modalities.
Extensive experiments demonstrate that a 3B-parameter model (Qwen2.5-VL-3B) trained under our framework outperforms comparable open-source baselines and advanced proprietary ones in multilingual visual reasoning. A further online A/B testing shows improvements in both answer acceptance rate and user satisfaction.
@@
  • We propose LaV-CoT, the first framework to unify language-aware visual CoT reasoning with multi-aspect reward optimization for multilingual text-centric VQA tasks.
@@
  • We design a language-aware GRPO algorithm with verifiable multi-aspect rewards, enhancing reasoning robustness and cross-modal alignment.
  • We validate our framework on public benchmarks and through an online A/B test, highlighting both state-of-the-art multilingual multimodal reasoning performance and its effectiveness in real-world deployment.
Figure 2. The framework includes an automated data generation pipeline, which leverages a multi-step reasoning process comprising (a) Text Summary with BBox, (b) Language Identification, (c) Spatial Object-level Image Captioning, and (d) Step-by-step Logical Reasoning. This reasoning pipeline first produces vanilla CoT annotations, which are then iteratively refined through rigorous verification to ensure high-quality supervision. The generated data subsequently supports a two-stage training paradigm combining SFT with GRPO. Multi-aspect Reward functions cover language consistency, count accuracy, the edit distance between predicted and ground-truth answers, and format compliance, collectively enhancing structural understanding and reinforcing robust reasoning capabilities.
Illustration of the multi-step data generation pipeline and two-stage training framework integrating SFT and GRPO for multilingual multimodal reasoning.
@@ 2. RELATED WORKS
2.1. Multilingual Text-Centric VQA in Vision-Language Models
Recent advances in vision-language models (VLMs) have significantly advanced visual question answering, particularly due to their strong zero-shot generalization capabilities. However, unlike conventional VQA, TEC-VQA represents a more complex challenge, particularly in multilingual settings, where cross-lingual text understanding and visual grounding must be performed jointly. Recent general models, such as Qwen2.5-VL (Bai et al., 2025) and InternVL3.5 (Wang et al., 2025), improve multilingual TEC-VQA performance by leveraging publicly available datasets spanning multiple languages. Expert models such as Parrot (Sun et al., 2025) further target multilingual text-centric reasoning by conditioning visual tokens on diverse language inputs and employing a Mixture-of-Experts (MoE) approach to enhance alignment across multilingual tokens.
2.2. Chain-of-Thought Reasoning in Vision Language Models
@@
CoT prompting (Wei et al., 2022) provides intermediate steps to tackle challenging problems such as commonsense reasoning and logical deduction (Cobbe et al., 2021). However, textual CoT methods rely solely on linguistic reasoning and often underutilize visual cues, resulting in weak visual grounding. Recent work, such as LLaVA-CoT (Xu et al., 2025), employs a structured CoT mechanism that systematically decomposes complex reasoning tasks using textual CoT enhanced with visual inputs, improving multimodal reasoning accuracy. Moreover, studies have shown that integrating CoT strategies significantly enhances the reasoning capability of VLMs, particularly in TEC-VQA. For instance, Visual CoT (Shao et al., 2024a) improves interpretability by generating bounding boxes that highlight relevant image regions alongside textual explanations, providing clearer visual grounding. Building upon these advancements, we propose a systematic and interpretable language-aware visual CoT pipeline specifically designed to address the challenges of multilingual vision-language reasoning.
2.3. Visual Reinforcement Fine-Tuning
With the emergence of reasoning models such as OpenAI’s o1 (OpenAI et al., 2024) and DeepSeek-R1 (Lu et al., 2023), research on large language models (LLMs) has increasingly focused on enhancing reasoning capabilities through reinforcement learning (RL). Recent studies extend this paradigm to vision-language models (VLMs) via RL-based post-training, known as Visual Reinforcement Fine-Tuning (Visual-RFT). For instance, Visual-RFT (Liu et al., 2025b) extends reinforcement fine-tuning to visual domains by training VLMs with multiple reasoning trajectories and optimizing them using visually verifiable rewards and policy gradient methods such as GRPO. Furthermore, Visionary-R1 (Xia et al., 2025) introduces a reinforcement learning framework that encourages visual understanding prior to reasoning by employing a structured caption–reason–answer training paradigm. In our work, we design verifiable multi-aspect reward functions to provide more granular and reliable feedback, enhancing reinforcement learning for multilingual TEC-VQA.
@@ 3. Method
As illustrated in Figure 2, our framework supports a progressive, step-by-step reasoning process that enhances the reasoning capabilities of VLMs. Building on the reasoning traces, we introduce an automated data curation pipeline that iteratively evaluates and refines each CoT sequence to ensure high-quality training data. Finally, a two-stage training strategy with a novel multi-aspect reward design is applied to produce the fully optimized model.
Dataset Language Size
COCO2017 EN,ZH,PT,AR,TR,RU 73k
Visual-Genome EN,ZH,PT,AR,TR,RU 18k
GQA EN,ZH,PT,AR,TR,RU 15k
OCR-VQA EN,ZH,PT,AR,TR,RU 16k
TextVQA EN,ZH,PT,AR,TR,RU 4.7k
Llava-Pretrain EN,ZH,PT,AR,TR,RU 0.6k
MTVQA AR,DE,FR,IT,JA,KO,RU,TH,VI 21k
3.1. Data Generation
3.1.1. Multi-stage Reasoning
The multi-stage reasoning design is motivated by the way humans naturally approach image understanding: when examining a document image, we first locate salient text regions and retain their summarized content, then identify the language, recognize objects and their spatial relationships, and finally integrate all information to reason step by step toward an answer. Here are the four structured stages:
@@ 3.1.1. Multi-stage Reasoning
3.1.2. Dataset Curation
Most existing multilingual VQA datasets lack the detailed reasoning processes necessary to effectively train the multilingual reasoning VLM model. To address this limitation, we compile a new dataset by integrating samples from several widely used VQA benchmarks, resulting in a total of 148k image–question–CoT–answer pairs, as illustrated in Table 1.
Specifically, as shown in Algorithm 1, we start from the original question–answer triplets $\langle I,Q,A\rangle$, where $I$ denotes the image, $Q$ the question, and $A$ the corresponding answer. We first prompt a generator $f_{\text{gen}}$ to produce an initial sequence of vanilla CoT steps $T_{\text{init}}=\{s_{i}\}_{i=1}^{|T_{\text{init}}|}$. Next, we prompt a evaluator $f_{\text{eval}}$ to score each step $s_{i}\in T_{\text{init}}$. Both $f_{\text{gen}}$ and $f_{\text{eval}}$ use GPT-4o-2024-05-13 (Hurst et al., 2024) with default decoding settings. For any step whose score falls below the threshold $\tau$, we iteratively perform the following procedure: First, we apply the evaluator to the step and then locate the erroneous part, denoted as $s_{\text{error}}$, using the function $\text{Locate}(f_{\text{eval}}(s_{i}))$. Next, a corrected step $s_{i}$ is generated by applying the function Correct to $s_{\text{error}}$, i.e., $s_{i}$ is updated as $\text{Correct}(f_{\text{gen}}(s_{\text{error}}))$. The corrected step then replaces the original step in the sequence $T_{\text{init}}$. Finally, the updated step is re-evaluated to obtain the new score using $f_{\text{eval}}(s_{i})$. This evaluation–correction–update loop repeats until all steps in $T_{\text{init}}$ exceed the threshold. The final verified Chain-of-Thought sequence is denoted as $T$, and the output dataset consists of $\langle I,Q,T,A\rangle$. Detailed prompt designs are shown in Appendix A.
Our dataset covers 13 languages, including English (EN), Chinese (ZH), Portuguese (PT), Arabic (AR), Turkish (TR), Russian (RU), German (DE), French (FR), Italian (IT), Japanese (JA), Korean (KO), Thai (TH), and Vietnamese (VI). These languages represent a diverse range of linguistic families. Our dataset is constructed from diverse sources, including COCO2017 (Lin et al., 2015), a large-scale dataset for object detection, segmentation, and image captioning containing everyday scenes with common objects in context; Visual Genome (Krishna et al., 2016), which provides densely annotated images linking objects, attributes, relationships, and region descriptions for visual reasoning; GQA (Hudson and Manning, 2019), emphasizing compositional reasoning over real-world images with complex question-answer pairs; OCR-VQA (Mishra et al., 2019) and TextVQA (Singh et al., 2019), both focusing on reading and reasoning over text within images for text-centric visual question answering; Llava-Pretrain (Liu et al., 2023a), a multimodal pretraining dataset supporting various VQA and reasoning tasks across multiple domains; and MTVQA (Tang et al., 2025), a multilingual text-based VQA dataset designed to evaluate cross-lingual multimodal understanding.